OpenStudy (shamil98):

Solving a system. x^2 + y^2 + 2x + 2y = 0 x^2 +y^2 + 4x + 6y + 12 = 0 hmm

OpenStudy (isaiah.feynman):

I really like the "hmmm". LOL

OpenStudy (shamil98):

Going about this algebraically will give me white hair and stress >.> and I must show all my work too .

OpenStudy (isaiah.feynman):

I have never solved this type of simultaneous equation.

OpenStudy (isaiah.feynman):

They look like equations of two circles.

OpenStudy (shamil98):

Yeah, they are circles.

OpenStudy (shamil98):

I know the two solutions but getting there is just :(

OpenStudy (isaiah.feynman):

Oh? what are x and y?

OpenStudy (shamil98):

One solution is (-2,-2)

OpenStudy (shamil98):

x^2 + y^2 + 2x + 2y = 0 y^2 = -x^2 - 2x - 2y x^2 + y^2 + 4x + 6y + 12 = 0 x^2 - x^2 - 2x - 2y + 4x + 6y + 12 =0 2x + 4y + 12 = 0 x + 2y = 6 was gonna do something like this but its still two variables

OpenStudy (shamil98):

x + 2y = -6*

OpenStudy (shamil98):

x = -2y - 6 (-2y-6)^2 + y^2 + 4(-2y-6) + 6y + 12 = 0 ?..

OpenStudy (isaiah.feynman):

I'm not sure, never solved circle equations this way!

OpenStudy (anonymous):

buddy why dont you substitute x^2+y^2 from 1st to 2nd then the same for 2nd to 1st u'll get 2 equations of x and y then its just simple

OpenStudy (shamil98):

Oh, why didnt i think of that LOL

OpenStudy (shamil98):

x^2 + y^2 = -2x -2y -2x -2y + 4x + 6y +12 = 0 2x + 4y + 12 = 0 -4x - 6y - 12 + 2x +2y = 0 so 2x + 4y = -12 -2x - 4y = 12 wait no that doesn't work, then its infinite solutions

OpenStudy (shamil98):

either that or i cant do algebra... prob the second one

OpenStudy (anonymous):

thats interesting .... :P then take eqn of y as x and put in 1s or 2nd one .......solve it as a quad and u'll get two values of x